OpenStudy (afloridagirl):

Solve 2x2 + 12x − 14 = 0 by completing the square. 2(x + 6)2 = 26; x = −6 + square root 13, x = −6 − square root 13 2(x + 6)2 = 20; x = −6 + square root 20, x = −6 − square root 20 2(x + 3)2 = 32; x = 1, x = −7 2(x + 3)2 = 32; x = 13, x = 19

OpenStudy (afloridagirl):

@jim_thompson5910 please help with this one

jimthompson5910 (jim_thompson5910):

\[\Large 2x^2+12x-14\] is the same as \[\Large 2x^2+12x+(-14)\]

jimthompson5910 (jim_thompson5910):

\[\Large {\color{red}{2}}x^2+{\color{green}{12}}x+({\color{blue}{-14}})\] is in the form \[\Large {\color{red}{a}}x^2+{\color{green}{b}}x+{\color{blue}{c}}\] where \[\Large \color{red}{a = 2}\] \[\Large \color{green}{b = 12}\] \[\Large \color{blue}{c = -14}\]

jimthompson5910 (jim_thompson5910):

making sense so far?

OpenStudy (afloridagirl):

i think so...

OpenStudy (afloridagirl):

would it be b?

jimthompson5910 (jim_thompson5910):

how did you get to b?

OpenStudy (afloridagirl):

im not sure tbh its kinda a guess

jimthompson5910 (jim_thompson5910):

ok let's forget about the answer choices for now. Let's just focus on each step by step

jimthompson5910 (jim_thompson5910):

do you see how I got \[\Large \color{red}{a = 2}\] \[\Large \color{green}{b = 12}\] \[\Large \color{blue}{c = -14}\]

OpenStudy (afloridagirl):

okay and we plug that in?

jimthompson5910 (jim_thompson5910):

plug a = 2 and b = 12 into h = -b/(2a) to get h = ???

OpenStudy (afloridagirl):

-3?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now we'll plug x = -3 back into y = 2x^2+12x-14 to get y = ???

OpenStudy (afloridagirl):

-32

jimthompson5910 (jim_thompson5910):

so the x coordinate of the vertex is -3, meaning h = -3 the y coordinate is -32, so k = -32

jimthompson5910 (jim_thompson5910):

plug in a = 2, h = -3, k = -32 into \[\Large y = a(x-h)^2 + k\] to get \[\Large y = 2(x-(-3))^2 + (-32)\] \[\Large y = 2(x+3)^2-32\]

OpenStudy (afloridagirl):

is it d?

jimthompson5910 (jim_thompson5910):

Plug in y = 0 and solve for x \[\Large y = 2(x+3)^2-32\] \[\Large 0 = 2(x+3)^2-32\] \[\Large 2(x+3)^2=32\] \[\Large (x+3)^2=16\] \[\Large x+3=4 \ \ \text{ or } \ \ x+3 = -4\] \[\Large x=1 \ \ \text{ or } \ \ x = -7\]

OpenStudy (afloridagirl):

oh so c?

OpenStudy (afloridagirl):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (afloridagirl):

thank you!

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (sshayer):

\[2x^2+12x-14=0,2\left( x^2+6x+\left( \frac{ 6 }{ 2 } \right)^2-\left( \frac{ 6 }{ 2 } \right)^2 \right)-14=0\] \[2\left( x+3 \right)^2-18-14=0,2(x+3)^2-32=0,(x+3)^2-16=0,(x+3)^2-4^2=0\] (x+3+4)(x+3-4)=0, (x+7)(x-1)=0 x=-7,1